OpenStudy (calculusxy):

Solve the equation 2x^3 - x^2 - 8x + 4 = 0 algebraically for all the values of x

OpenStudy (calculusxy):

@Kainui

Directrix (directrix):

Try factoring by grouping to get at least one root. 2x^3 - x^2 - 8x + 4 = 0 = x^2 ( 2x - 1) - 4 ( 2x - 1) =

OpenStudy (calculusxy):

Okay. And then?

Directrix (directrix):

x^2 ( 2x - 1) - 4 ( 2x - 1) = The common factor is (2x - 1)

OpenStudy (calculusxy):

ok

Directrix (directrix):

x^2 ( 2x - 1) - 4 ( 2x - 1) = ( 2x - 1) ( x^2 - 4)

Directrix (directrix):

( 2x - 1) ( x^2 - 4) - 0 Use the Zero Product Property to solve for x. Note that x^2 - 4 is the difference of 2 squares.

OpenStudy (calculusxy):

ok gimme a moment

Directrix (directrix):

Alrighty. Take your time. "Fast is fine, but accuracy is everything." ~Wyatt Earp

OpenStudy (calculusxy):

so i have (x+2)(x-2)(2x-1) x = -2, 2, 0

Directrix (directrix):

How did you get 0 for this: (2x-1)= 0

OpenStudy (calculusxy):

oh sorry. i meant 1/2

Directrix (directrix):

1/2, 2, and -2 Correct.

OpenStudy (calculusxy):

i have other questions can i post them for help?

OpenStudy (calculusxy):

and thanks for ur help

Directrix (directrix):

Yes. Start a new thread, okay?

OpenStudy (calculusxy):

ok

Directrix (directrix):

You are welcome.